OpenStudy (chillout):

Another linear algebra question... Proof of: If A and B are nonsingular matrices, (A+B) is not necessarily invertible.

OpenStudy (chillout):

I'm just assuming A = I and B=-I. Is there a generalization for this?

OpenStudy (anonymous):

you can do this proof by example think up two matrices whose determinant is not zero, but the sum has determinant zero examples abound

OpenStudy (vishweshshrimali5):

Well for proving negation if you can provide even one case then you are good to go

OpenStudy (chillout):

I used an example to prove. But is there any formal way to describe this proof?

OpenStudy (vishweshshrimali5):

Well you can if you want...

OpenStudy (vishweshshrimali5):

Let's assume that for every non singular and invertible A and B, A+B is non singular and invertible... Then that means... \(|A| \ne 0\) and \(|B| \ne 0\) Now, \(|A+B| = |A| + |B| \ne 0 \iff |A| \ne -|B|\) But this is possible... you can here provide the example of A = I and B = -I

OpenStudy (chillout):

I thought as much... All the proofs I'm coming along are long but this one seems simple enough.

OpenStudy (chillout):

Using \(det(A)\) method is out of scope for now, though

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (chillout):

Seems I can't help but use it xD. All right, thanks for the help again!

OpenStudy (vishweshshrimali5):

:) Glad to be of help

OpenStudy (chillout):

Any other proofs?

OpenStudy (chillout):

Not using determinants, I mean. For now it's out of scope;

OpenStudy (kainui):

I was only gonna say this, but it's so silly I decided not to. Oh well here it is haha If A is nonsingular, B=-A is nonsingular but clearly A+B = A-A = 0 is singular. So it's a mild generalization of what you were asking but nothing too special really haha.

OpenStudy (chillout):

All right, fair enough xD

OpenStudy (chillout):

So a general approach is A+B=0. Not too much mathematical formalism but I'm OK with this XD

OpenStudy (kainui):

There's probably better answers out there, but this is pretty good to me haha.

OpenStudy (chillout):

All right, thanks!